OpenStudy (anonymous):

Evaluate d/dx of x^x^2

OpenStudy (anonymous):

y=x^x^2 lny=x^2(lnx) Differentiate this implicitly y'/y=2xlnx+x y'=y(2xlnx+x) y'=x^x^2(2xlnx+x)

OpenStudy (anonymous):

Great, now... how would you solve 0=x^x^2(2xlnx+x) ... I know it is 1/sqrt(e).. I just need help with the intermediate steps